OpenStudy (anonymous):

Graph the function. Then use your graph to find the indicated limit. f(x) = 7ex , lim x -> 0 f(x) A. 0 B. 7 C. 1 D. -7

OpenStudy (raden):

you mean : 

OpenStudy (anonymous):

YES

OpenStudy (raden):

ok, because x->0, just subtitute x =0 to f(x) it is 7 e^0 = 7 * 1 = 7

OpenStudy (anonymous):

OK THANK YOU @RadEn

OpenStudy (raden):

welcome